OpenStudy (anonymous):

-6p-21=3p-12 help me check my answer im not sure if is right

OpenStudy (anonymous):

wats ur answer?

OpenStudy (anonymous):

p=-1

OpenStudy (anonymous):

-6p-21=3p-12 -6p-21+12=3p-12+12 -6p-9=3p -6p+6p-9=3p+6p -9=9p -9/9=9p/9 -1=p p=-1